MsBrains (ms-brains):

PLEASE HELP! I ONLY HAVE 2 QUESTIONS! Can someone please help me fill out this Practice sheet for Algebra1!! I WILL MEDAL AND FAN! CX

MsBrains (ms-brains):

Try to follow the Rubric on the side

OpenStudy (qwertty123):

Having a hard time seeing it!

OpenStudy (qwertty123):

What do you really need help with?

MsBrains (ms-brains):

You clicked on the link, then you can zoom in

OpenStudy (study312):

Let me see if this size is ok

OpenStudy (study312):

If you click on that pic I just posted you can click again to make it bigger

OpenStudy (study312):

Is that good @Qwertty123?

MsBrains (ms-brains):

@Study312 or @Qwertty123 Could one of you please help?

OpenStudy (study312):

Sorry Ms-Brains this is not my area of expertise. I have not quite learned this yet.

OpenStudy (qwertty123):

Oh hi sorry sure Let me see the question again!

OpenStudy (qwertty123):

Still need help @Study312

OpenStudy (qwertty123):

I mean @Ms-Brains

MsBrains (ms-brains):

Yes!

OpenStudy (qwertty123):

okay just wait!

OpenStudy (qwertty123):

Can you solve it?

OpenStudy (qwertty123):

The line will be going down!

OpenStudy (qwertty123):

Hello @Ms-Brains

MsBrains (ms-brains):

I have no idea how to solve :(

MsBrains (ms-brains):

Also, I will be right back! Give me 5mins!

OpenStudy (qwertty123):

oh okay well here is my answer I got! y≤−x/3+2!

OpenStudy (qwertty123):

The line is going down \ and you would shade down from the line! \ Down! I have to go! Sorry @Ms-Brains ! Hopefully some one else can explain for you! :)

OpenStudy (anonymous):

consider x+3y=6 first we graph it. when x=0,3y=6,y=2 so point A is (0,2) when y=0,x=6 so point B is (6,0) put x=0,y=0 in \[x+3y \le 6,0+0\le 6,0\le6 \] which is true. Now we make agraph 

MsBrains (ms-brains):

Ok